OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values.

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

so we know this is a two-step function right?

OpenStudy (texaschic101):

try subbing in a value of x to find y.... lets let x = 2 now we sub y = -2x - 3 y = -2(2) - 3 y = -4 - 3 y = -7 so when x = 2, then y = -7 you can continue doing it this way and you will have your table of values.